You’ve decided to set aside $5,000 in a commercial bank certificate of deposit (CD). Assuming your money will earn interest at a rate of 3%, how much will your entire account be worth at the end of the first year?

Group of answer choices

$15,000

$5,150

$150

$1,666
Answer #1

Answer

Interest amount =investment * interest rate =5000*0.03=150

total amount =investment +interest earned =5000+150=5150

your entire account will be $5150

Option 2